The owner of a health club franchise believes that the average age of the clubs members is under 45 years. To test this, a rHello, I need some help with the second question?

Thank you!
Answer #1

Sample size = n = 30

Degree of freedom =n-1 = 30 - 1 = 29

Left tailed test.

t critical value is = -1.70

Eth $5000 ਲਾ ਦੇ ਨਵ ਕਿ ਜੋ % % ਉਸ * t Table * *